# Quantum information with quantum-like bits

### v5

@@ II.2.2 Example: Bell states
$\mathcal{R}=\mathcal{R}^{(1)}\otimes\mathbb{1}_{2{N_{G}}}+\mathbb{1}_{2{N_{G}}}\otimes\mathcal{R}^{(2)}.$ (II.19)
@@
$\rho(\lambda)=\int_{-\infty}^{+\infty}\mathrm{d}x\;\rho^{(1)}(x)\rho^{(2)}(\lambda-x),$ (II.20)
@@
$\psi^{(1)}_{\uparrow/\downarrow}=\begin{pmatrix}a_{1}\\ \pm a_{2}\end{pmatrix},\hskip 14.22636pt\psi^{(2)}_{\uparrow/\downarrow}=\begin{pmatrix}b_{1}\\ \pm b_{2}\end{pmatrix},$ (II.21)
where $a_{i}$’s and $b_{i}$’s are defined as in ?? for the resources ${\mathcal{R}}^{(1)}$ and ${\mathcal{R}}^{(2)}$, respectively. The notation in ?? indicate that the upper arrow ($\uparrow$) on the left-hand side is associated to the positive sign on right-hand side, and similarly for $\downarrow$ and the negative sign. With ?? we can express
$\psi_{\uparrow,\uparrow/\downarrow}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}\\ \pm a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}\\ \pm a_{2}\otimes b_{2}\end{pmatrix},\hskip 14.22636pt\psi_{\downarrow,\uparrow\downarrow}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}\\ \pm a_{1}\otimes b_{2}\\ -a_{2}\otimes b_{1}\\ \mp a_{2}\otimes b_{2}\end{pmatrix}.$ (II.22a)
@@
$\Phi^{\mathrm{Bell}}_{\uparrow/\downarrow}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{\uparrow\uparrow}\pm\psi_{\downarrow\downarrow}\right),$ (II.23a)
$\Psi^{\mathrm{Bell}}_{\uparrow/\downarrow}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{\uparrow\downarrow}\pm\psi_{\downarrow\uparrow}\right),$ (II.23b)
@@
$\Phi^{\mathrm{Bell}}_{\uparrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}\\ \mathbb{0}_{{N_{G}}}\\ \mathbb{0}_{{N_{G}}}\\ a_{2}\otimes b_{2}\end{pmatrix},$ $\Phi^{\mathrm{Bell}}_{\downarrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\mathbb{0}_{{N_{G}}}\\ a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}\\ \mathbb{0}_{{N_{G}}}\end{pmatrix},$ (II.24a)
$\Psi^{\mathrm{Bell}}_{\uparrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}\\ \mathbb{0}_{{N_{G}}}\\ \mathbb{0}_{{N_{G}}}\\ -a_{2}\otimes b_{2}\end{pmatrix},$ $\Psi^{\mathrm{Bell}}_{\downarrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\mathbb{0}_{{N_{G}}}\\ -a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}\\ \mathbb{0}_{{N_{G}}}\end{pmatrix}.$ (II.24b)
@@ III Classical synchronization
Here, we investigate classical synchronization in dynamical systems of oscillators, where two-body correlations are governed by a connectivity matrix with the structure described in ??. Systems that align with our formalism must, as a general requirement, involve classical many-body dynamics, where the two-body couplings between degrees of freedom are governed by a connectivity matrix $\mathcal{R}$, as outlined in ??. Additionally, the dynamics should guarantee that synchronization leads to a globally stable state that remains robust over long time scales. These features are satisfied by the non-symplectic dynamics of the Kuramoto model Kuramoto (1975); Kawamura et al. (2010); Scholes (2022); Rodrigues et al. (2016), which is characterized by its nonlinear equations of motion
$\dot{\theta}_{i}=\omega_{i}+\frac{\Gamma}{2{N_{G}}}\sum_{j=1}^{2{N_{G}}}\mathcal{R}_{ij}\sin(\theta_{j}-\theta_{i}),\hskip 19.91692pti=1,\cdots,2{N_{G}}.$ (III.1)
Here, the $\theta_{i}$’s are angular variables corresponding to $2{N_{G}}$ oscillators, defined on the vertices of a graph. The matrix $\mathcal{R}$ encodes the correlations between these degrees of freedom. The $\omega_{i}$’s denote the intrinsic oscillation frequencies. The time evolution of the angular variables of the model is shown in panel a) of ??. For our chosen system parameters (see caption of the figure) the system reaches a synchronized steady state, where all oscillators evolve at a uniform, constant phase. The synchronized steady state appears robust even in presence of disorder in the network, as shown in panel b) of the same figure. Here, white static disorder, with variance $\sigma=3$ and a zero mean, is added to all nonzero entries of the adjacency matrix $\mathcal{R}$, originally valued at one. Despite this strong perturbation, a collective synchronization state persists at long times, with only minor phase shifts affecting small subgroups of oscillators. This observation further supports the authors’ previous findings on the resilience of synchronization dynamics to noise–a property particularly appealing for robust information encoding and processing in the steady states of these classical dynamical systems. Scholes and Amati (2025); Scholes (2023, 2024)
@@
$\mathcal{R}=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\lambda_{\bm{\sigma}}\left(\psi_{\bm{\sigma}}\otimes\psi_{\bm{\sigma}}^{\dagger}\right)+\sum_{i=1}^{M-2}\tilde{\lambda}_{i}\left(\tilde{\psi}_{i}\otimes\tilde{\psi}_{i}^{\dagger}\right),$ (III.2)
where $M=(2{N_{G}})^{N_{\mathrm{QL}}}$ denotes the total number of states. The first term in ?? represents the contribution of the emergent states $\psi_{\bm{\sigma}}$’s. The second term corresponds to the random part of the spectrum, with eigenvalues $\tilde{\lambda}_{i}$’s and corresponding eigenvectors $\tilde{\psi}_{i}$’s. Panel b) of ?? depicts the time evolution of the Kuramoto dynamics when only the random part of the spectrum is taken into account [that is, by imposing $\lambda_{\bm{\sigma}}\equiv 0$ $\forall\bm{\sigma}\in\bm{\mathcal{S}}$ in ??]. In this scenario, synchronization fails to emerge, with all degrees of freedom maintaining their initial random phases throughout the entire time evolution. This highlights that the structure of two-body correlations and emergent states, encoded in the adjacency matrix $\mathcal{R}$, play a critical role in achieving classical synchronization. In the next ??, we build on these idea by exploring the connection between dynamical synchronization and information processing. This approach provides a framework for linking the steady states of classical many-body systems to quantum information protocols.
Figure 5: a): Time evolution of a system with $602{N_{G}}=$ 60 oscillators governed by the Kuramoto model [??]. The angular variables $\theta_{i}$ are initialized uniformly in $[0,2\pi)$, and the intrinsic frequencies $\omega_{i}$ are sampled uniformly from $(0,1)$. The coupling constant is set to $\Gamma=10$, and the connectivity matrix $\mathcal{R}$ is structured as a regular graph, as described in ??, with the valencies of the diagonal and off-diagonal blocks set to $k=29$ and $l=15$, respectively. b): Time evolution of the system under the same parameters and initial conditions as in a), but with disorder introduced into the connectivity matrix $\mathcal{R}$. Specifically, each nonzero entry of $\mathcal{R}$ is perturbed by random white noise with zero mean and variance $\sigma=3$. c): Dynamics of the Kuramoto model with the connectivity matrix restricted to its random components [see discussion following ??]. The system parameters and initial conditions are the same as in a) and b).
@@ IV.1 Formalism
To connect the usual definition of gates acting on an abstract Hilbert space to our QL formalism, it is convenient to introduce a map $U_{\mathrm{cb}}$ transforming the emergent eigenvectors onto the standard computational basis. As that all elementary subrgaphs in our resources [????] are regular, this map can be expressed analytically by
@@
$U_{\mathrm{cb}}\psi_{\uparrow}=\begin{pmatrix}1_{N_{G}}\\ 0_{N_{G}}\end{pmatrix},\hskip 14.22636ptU_{\mathrm{cb}}\psi_{\downarrow}=\begin{pmatrix}0_{N_{G}}\\ 1_{N_{G}}\end{pmatrix},$ (IV.3)
and
$x_{M}=(x,\cdots,x)/\sqrt{M}\in\mathbb{C}^{M},\hskip 28.45274ptx\in\mathbb{C},\;M\in\mathbb{N}.$ (IV.4)
Note that, given that all the basis subgraphs are regular, the solution to ?? is independent of their local structure. We show in ?? that a solution for the map $U_{\mathrm{cb}}$ can be determined if the resources are not exactly regular, by solving numerically a linear algebra problem. We define the QL analog of any arbitrary quantum gate $V_{\mathrm{g}}$ by the matrix transformation
$U_{\mathrm{g}}=U_{\mathrm{cb}}^{-1}(V_{\mathrm{g}}\otimes\mathbb{1}_{{N_{G}}})U_{\mathrm{cb}}.$ (IV.5)
@@
$U_{\mathcal{I}}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\end{pmatrix},$ (IV.6a)
$U_{X}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&-\mathbb{1}_{N_{G}}\end{pmatrix},$ (IV.6b)
$U_{\mathrm{Y}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&i\mathbb{1}_{N_{G}}\\ -i\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (IV.6c)
$U_{\mathrm{Z}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\\ \mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (IV.6d)
respectively. As expected, the transformations in ?? act as the standard Pauli gates:
@@
$U_{\mathrm{H}}\psi_{\uparrow/\downarrow}=\left(V_{\mathrm{H}}\otimes\mathbb{1}_{{N_{G}}}\right)\psi_{\uparrow/\downarrow}=\frac{1}{\sqrt{2}}(\psi_{\uparrow}\pm\psi_{\downarrow}).$ (IV.8)
The formal approach proposed here applies equally to any arbitrary gates, including those outside the Clifford group. While gates outside the Clifford group break the stabilizer structure and require exponentially more resources compared to Clifford gates—which can be efficiently simulated classically—the distinction between these two classes of gates is not relevant to our framework. Our formalism is designed to be computationally exact, addressing the exponentially high cost associated with scaling QL resources, regardless of the type of gate involved. The action of the non-Clifford T-gate on the factorized basis is
$U_{\mathrm{T}}\psi_{\uparrow/\downarrow}=\frac{1}{2}\left[(1+\,\mathrm{e}^{i\pi/4})\psi_{\uparrow}+(1-\,\mathrm{e}^{i\pi/4})\psi_{\downarrow}\right].$ (IV.9)
[…]
$U_{\mathrm{CN}}$ $=\frac{1}{2}(U_{\mathcal{I}}+U_{\mathrm{Z}})\otimes U_{\mathcal{I}}+\frac{1}{2}(U_{\mathcal{I}}-U_{\mathrm{Z}})\otimes U_{\mathrm{X}}$ (IV.10)
@@
$U_{\mathrm{CN}}\psi_{\uparrow[\uparrow,\downarrow]}=\psi_{\uparrow[\uparrow,\downarrow]},\hskip 28.45274ptU_{\mathrm{CN}}\psi_{\downarrow[\uparrow,\downarrow]}=\psi_{\downarrow[\downarrow,\uparrow]}.$ (IV.11)
Up to this point, we have focused exclusively on the action of QL gates on states $\psi_{\bm{\sigma}}$ of the factorized basis. We can infer from those the transformation of any arbitrary state $\psi=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\gamma_{\bm{\sigma}}\psi_{\bm{\sigma}}$, $\gamma_{\bm{\sigma}}\in\mathbb{C}$, thanks to linearity of the map ??. From a practical standpoint, it is preferable to translate these state operations onto transformations applicable directed to the graph resources. This feature is a key requirement for a future experimental validation of our formalism with experimental resources. (An extended discussion on experimental resources is provided in the next ??). A straightforward approach in this regard is to define the unitary map
$\mathcal{R}\mapsto\mathcal{R}_{\mathrm{g}}=U_{\mathrm{g}}\mathcal{R}U_{\mathrm{g}}^{\dagger}.$ (IV.12)
?? transforms all eigenstates of a resource consistently to the gate $U_{\mathrm{g}}$ as defined in ??. This property can be easily inferred by expanding $\mathcal{R}$ on the right-hand side of ?? via the spectral decomposition in ??. In the next ?? we provide concrete examples of ??, and we discuss how this network transformation affects the steady-state synchronizing solution of a system of oscillators.

### v6

@@ II.2.2 Example: Bell states
$\mathcal{R}=\mathcal{R}^{(1)}\otimes\mathbb{1}_{2{N_{G}}}+\mathbb{1}_{2{N_{G}}}\otimes\mathcal{R}^{(2)}.$ (II.14)
@@
$\rho(\lambda)=\int_{-\infty}^{+\infty}\mathrm{d}x\;\rho^{(1)}(x)\rho^{(2)}(\lambda-x),$ (II.15)
@@
$\psi^{(1)}_{\uparrow/\downarrow}=\begin{pmatrix}a_{1}\\ \pm a_{2}\end{pmatrix},\hskip 14.22636pt\psi^{(2)}_{\uparrow/\downarrow}=\begin{pmatrix}b_{1}\\ \pm b_{2}\end{pmatrix},$ (II.16)
where $a_{i}$’s and $b_{i}$’s are defined as in ?? for the resources ${\mathcal{R}}^{(1)}$ and ${\mathcal{R}}^{(2)}$, respectively. The notation in ?? indicate that the arrow on the left (right) of the slash is associated to the upper (lower) sign on the right-hands side. With ?? we can express
$\psi_{\uparrow,\uparrow/\downarrow}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}\\ \pm a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}\\ \pm a_{2}\otimes b_{2}\end{pmatrix},\hskip 14.22636pt\psi_{\downarrow,\uparrow/\downarrow}$ $=\frac{1}{2}\begin{pmatrix}a_{1}\otimes b_{1}\\ \pm a_{1}\otimes b_{2}\\ -a_{2}\otimes b_{1}\\ \mp a_{2}\otimes b_{2}\end{pmatrix}.$ (II.17a)
@@
$\Phi^{\mathrm{Bell}}_{\uparrow/\downarrow}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{\uparrow\uparrow}\pm\psi_{\downarrow\downarrow}\right),$ (II.18a)
$\Psi^{\mathrm{Bell}}_{\uparrow/\downarrow}$ $=\tfrac{1}{\sqrt{2}}\left(\psi_{\uparrow\downarrow}\pm\psi_{\downarrow\uparrow}\right),$ (II.18b)
@@
$\Phi^{\mathrm{Bell}}_{\uparrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}\\ 0_{N_{\mathrm{G}}^{2}}\\ 0_{N_{\mathrm{G}}^{2}}\\ a_{2}\otimes b_{2}\end{pmatrix},$ $\Phi^{\mathrm{Bell}}_{\downarrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}0_{N_{\mathrm{G}}^{2}}\\ a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}\\ 0_{N_{\mathrm{G}}^{2}}\end{pmatrix},$ (II.19a)
$\Psi^{\mathrm{Bell}}_{\uparrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}a_{1}\otimes b_{1}\\ 0_{N_{\mathrm{G}}^{2}}\\ 0_{N_{\mathrm{G}}^{2}}\\ -a_{2}\otimes b_{2}\end{pmatrix},$ $\Psi^{\mathrm{Bell}}_{\downarrow}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}0_{N_{\mathrm{G}}^{2}}\\ -a_{1}\otimes b_{2}\\ a_{2}\otimes b_{1}\\ 0_{N_{\mathrm{G}}^{2}}\end{pmatrix},$ (II.19b)
where
$x_{M}=(x,\cdots,x)/\sqrt{M}\in\mathbb{C}^{M},\hskip 28.45274ptx\in\mathbb{C},\;M\in\mathbb{N}.$ (II.20)
@@ III Classical synchronization
Here, we investigate classical synchronization in dynamical systems of oscillators, where two-body correlations are governed by a connectivity matrix with the structure described in ??. Systems that align with our formalism must, as a general requirement, involve classical many-body dynamics, where the two-body couplings between degrees of freedom are governed by a connectivity matrix $\mathcal{R}$, as outlined in ??. Additionally, the dynamics should guarantee that synchronization leads to a globally stable state that remains robust over long time scales. These features are satisfied by the Kuramoto model, Kuramoto (1975); Kawamura et al. (2010); Scholes (2022); Rodrigues et al. (2016) described by the nonlinear and non-symplectic equations of motion
$\dot{\theta}_{i}=\omega_{i}+\frac{\Gamma}{2{N_{G}}}\sum_{j=1}^{2{N_{G}}}\mathcal{R}_{ij}\sin(\theta_{j}-\theta_{i}),\hskip 19.91692pti=1,\ldots,2{N_{G}}.$ (III.1)
Here, the $\theta_{i}$’s are angular variables corresponding to $2{N_{G}}$ oscillators, defined on the vertices of a graph. The matrix $\mathcal{R}$ encodes the correlations between these degrees of freedom. The $\omega_{i}$’s denote the intrinsic oscillation frequencies. The time evolution of the angular variables of the model is shown in panel a) of ??. For our chosen system parameters (see caption of the figure) the system reaches a synchronized steady state, where all oscillators evolve at a uniform, constant phase. The synchronized steady state appears robust even in presence of disorder in the network, as shown in panel b) of the same figure. Here, white static disorder, with variance $\sigma=3$ and a zero mean, is added to all nonzero entries of the adjacency matrix $\mathcal{R}$, originally valued at one. Despite this strong perturbation, a collective synchronization state persists at long times, with only minor phase shifts affecting small subgroups of oscillators. This observation further supports the authors’ previous findings on the resilience of synchronization dynamics to noise—a property particularly appealing for robust information encoding and processing in the steady states of these classical dynamical systems. Scholes and Amati (2025); Scholes (2023, 2024)
@@
$\mathcal{R}=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\lambda_{\bm{\sigma}}\psi_{\bm{\sigma}}\psi_{\bm{\sigma}}^{\dagger}+\sum_{i=1}^{M}\tilde{\lambda}_{i}\tilde{\psi}_{i}\tilde{\psi}_{i}^{\dagger},$ (III.2)
where $M=(2{N_{G}})^{N_{\mathrm{QL}}}-2^{N_{\mathrm{QL}}}$. The first term in ?? represents the contribution of the emergent states $\psi_{\bm{\sigma}}$’s. The second term corresponds to the random part of the spectrum, with eigenvalues $\tilde{\lambda}_{i}$’s and corresponding eigenvectors $\tilde{\psi}_{i}$’s. In panel b) of ??, we show the time evolution of a system initialized with the same conditions as in panel a) and governed by the same interaction rules, with the only difference being the addition of static random noise $n$. This noise has zero mean, variance $\sigma_{\mathrm{n}}=3$, and is sampled independently for each nonzero component of the adjacency matrix. Despite the noise being strong compared to the oscillator coupling strength (set to 1), the system still shows a clear synchronization pattern over time. This provides numerical evidence that, under broad conditions, the dynamics of the model can be highly resilient to static disorder. Such robustness is related to the collective behavior that emerges through synchronization. Finally, we show in panel c) of ?? the time evolution of the Kuramoto dynamics when only the random part of the spectrum is taken into account [that is, by imposing $\lambda_{\bm{\sigma}}\equiv 0$ $\forall\bm{\sigma}\in\bm{\mathcal{S}}$ in ??]. In this scenario, synchronization fails to emerge, with all degrees of freedom maintaining their initial random phases throughout the entire time evolution. This highlights that the structure of two-body correlations and emergent states, encoded in the adjacency matrix $\mathcal{R}$, play a critical role in achieving classical synchronization. In the next ??, we build on these idea by exploring the connection between dynamical synchronization and information processing. This approach provides a framework for linking the steady states of classical many-body systems to quantum information protocols.
Figure 5: a): Time evolution of a system with $602{N_{G}}=$ 60 oscillators governed by the Kuramoto model [??]. The angular variables $\theta_{i}$ are initialized uniformly in $[0,2\pi)$, and the intrinsic frequencies $\omega_{i}$ are sampled uniformly from $[0,1]$. The coupling constant is set to $\Gamma=10$, and the connectivity matrix $\mathcal{R}$ is structured as a regular graph, as described in ??, with the valencies of the diagonal and off-diagonal blocks set to $k=29$ and $l=15$, respectively. b): Time evolution of the system under the same parameters and initial conditions as in a), but with disorder introduced into the connectivity matrix $\mathcal{R}$. Specifically, each nonzero entry of $\mathcal{R}$ is perturbed by random white noise with zero mean and variance $\sigma_{\mathrm{n}}=3$. c): Dynamics of the Kuramoto model with the connectivity matrix restricted to its random components [see discussion following ??]. The system parameters and initial conditions are the same as in a) and b).
@@ IV.1 Formalism
To connect the usual definition of gates acting on an abstract Hilbert space to our QL formalism, it is convenient to introduce a map $U_{\mathrm{cb}}$ transforming the emergent eigenvectors onto the standard computational basis. Given that all elementary subrgaphs in our resources [????] are regular, this map can be expressed analytically by
@@
$U_{\mathrm{cb}}\psi_{\uparrow}=\begin{pmatrix}1_{N_{G}}\\ 0_{N_{G}}\end{pmatrix},\hskip 14.22636ptU_{\mathrm{cb}}\psi_{\downarrow}=\begin{pmatrix}0_{N_{G}}\\ 1_{N_{G}}\end{pmatrix}.$ (IV.3)
Note that, given that all the basis subgraphs are regular, the solution to ?? is independent of their local structure. We show in ?? that a solution for the map $U_{\mathrm{cb}}$ can be determined even if the resources are not exactly regular, by solving numerically a linear algebra problem. We define the QL analog of any arbitrary quantum gate $V_{\mathrm{g}}$ by the matrix transformation
$U_{\mathrm{g}}=U_{\mathrm{cb}}^{-1}(V_{\mathrm{g}}\otimes\mathbb{1}_{{N_{G}}})U_{\mathrm{cb}}.$ (IV.4)
@@
$U_{\mathcal{I}}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\end{pmatrix},$ (IV.5a)
$U_{\mathrm{X}}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&-\mathbb{1}_{N_{G}}\end{pmatrix},$ (IV.5b)
$U_{\mathrm{Y}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&i\mathbb{1}_{N_{G}}\\ -i\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (IV.5c)
$U_{\mathrm{Z}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\\ \mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (IV.5d)
respectively. As expected, the transformations in ?? act as the standard Pauli gates, namely,
$U_{\mathrm{X}}\psi_{\uparrow/\downarrow}=\psi_{\downarrow/\uparrow},\hskip 14.22636ptU_{\mathrm{Y}}\psi_{\uparrow/\downarrow}=\pm i\psi_{\downarrow/\uparrow},\hskip 14.22636ptU_{\mathrm{Z}}\psi_{\uparrow/\downarrow}=\pm\psi_{\uparrow/\downarrow}.$ (IV.6)
$U_{\mathrm{H}}\psi_{\uparrow/\downarrow}=\left(V_{\mathrm{H}}\otimes\mathbb{1}_{{N_{G}}}\right)\psi_{\uparrow/\downarrow}=\frac{1}{\sqrt{2}}(\psi_{\uparrow}\pm\psi_{\downarrow}).$ (IV.7)
The approach proposed here applies equally to any kind of gate, including those outside the Clifford group. This class of gates break the stabilizer structure, making them harder to simulate compared to Clifford gates. In contrast, Clifford gates can be simulated efficiently on a classical computer. However, this difference between Clifford and non-Clifford gates does not matter in our framework. Our formalism already takes into account the exponential cost of scaling QL resources, no matter what type of gate is used. The action of the non-Clifford T-gate on the factorized basis is
$U_{\mathrm{T}}\psi_{\sigma}=\left(\delta_{\sigma,\uparrow}+\,\mathrm{e}^{i\pi/4}\delta_{\sigma,\downarrow}\right)\psi_{\sigma}.$ (IV.8)
@@
$U_{\mathrm{CN}}$ $=\frac{1}{2}(U_{\mathcal{I}}+U_{\mathrm{Z}})\otimes U_{\mathcal{I}}+\frac{1}{2}(U_{\mathcal{I}}-U_{\mathrm{Z}})\otimes U_{\mathrm{X}}$ (IV.9)
@@
$U_{\mathrm{CN}}\psi_{\uparrow,\uparrow/\downarrow}=\psi_{\uparrow,\uparrow/\downarrow},\hskip 28.45274ptU_{\mathrm{CN}}\psi_{\downarrow,\uparrow/\downarrow}=\psi_{\downarrow,\downarrow/\uparrow}.$ (IV.10)
Up to this point, we have focused exclusively on the action of QL gates on states $\psi_{\bm{\sigma}}$ of the factorized basis. We can infer from those the transformation of any arbitrary state $\psi=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\gamma_{\bm{\sigma}}\psi_{\bm{\sigma}}$, $\gamma_{\bm{\sigma}}\in\mathbb{C}$ from the linearity of the map ??. From a practical standpoint, it is preferable to translate these state operations onto transformations applicable directed to the graph resources. This feature is a key requirement for a future experimental validation of our formalism with experimental resources. (An extended discussion on experimental resources is provided in the next ??). A straightforward approach in this regard is to define the unitary map
$\mathcal{R}\mapsto\mathcal{R}_{\mathrm{g}}=U_{\mathrm{g}}\mathcal{R}U_{\mathrm{g}}^{\dagger}.$ (IV.11)
?? transforms all eigenstates of a resource consistently to the gate $U_{\mathrm{g}}$ as defined in ??. This can be easily inferred by expanding $\mathcal{R}$ on the right-hand side of ?? via the spectral decomposition in ??. This map can be straightforwardly extended to arbitrary quantum circuits, described by the sequence of $M$ gate operations $\bm{\mathrm{g}}=\{\mathrm{g}_{1},\ldots,\mathrm{g}_{M}\}$. In this case, we define by
$U_{\bm{\mathrm{g}}}=U_{\mathrm{cb}}^{-1}(V_{\bm{\mathrm{g}}}\otimes\mathbb{1}_{{N_{G}}})U_{\mathrm{cb}}=U_{\bm{\mathrm{g}}}=U_{\mathrm{g}_{M}}\cdots U_{\mathrm{g}_{1}},$ (IV.12)
the corresponding unitary map acting on QL states, with an analog network transformation defined by
$\mathcal{R}_{\bm{\mathrm{g}}}=U_{\bm{\mathrm{g}}}\mathcal{R}U_{\bm{\mathrm{g}}}^{\dagger}.$ (IV.13)
In the next ?? we provide concrete examples of ??, and we discuss how this network transformation affects the steady-state synchronizing solution of a system of oscillators. A more general analysis of circuit transformations, associated to ??, can be found in Ref. Amati and Scholes, 2025.